+ +

IV Journal Structure

1 Text Formats

A typeset layout such as this one displays the structure of a text directly to a human reader. Many different electronic data formats can produce the same layout.

Typeset Layout of a Sample Text

+ +

[figure]

Geometric Data Format

+ +

page

size

3 in. × 3 in.

margins

t: .5 in, b: .5 in, l: .25 in, r: .25 in

contents

line

type

12pt boldface

justify

justify left, ragged right

contents

IV Journal Structure

space

.375 in

contents

line

type

11pt boldface

justify

justify left, ragged right

contents

1 Text Formats

space

.25 in

contents

line

type

10pt roman

justify

justify left and right

contents

A typeset layout such
as this one dis-

⋮

Typographic Format

+ +

[figure: I; V; J; o; u; r; n; a; l]

Sequential Format

+ +

[figure: section; number; 4; title; Journal Structure; contents; subsection; number; 1; title; Text Formats; contents; paragraph; contents; sentence; contents; A typeset ; layout such ; as this ⋯; ⋮; ⋮]

Structural Format

+ +

[figure: normal function; actor; typeset layout; role; communication; medium; visual; receiver; human]

Semantic Format

+ +

Summary of Textual Formats

1. Geometric • Bitmap, Pixel Image

• dots, coordinates

2. Typographic • PostScript, DVI, MacWrite, TEX, troff

• characters, fonts, point sizes, lines, pages

3. Sequential • ASCII, Character Set

• characters, cases, control characters

4. Structural • None yet—LaTEX, AMSTEX, SGML are close • sections, paragraphs, titles, sentences

5. Semantic • Knowledge Base, Semantic Net

• concepts, objects, predicates, relations

+ +

5 Roles in Scholarly Publication

1. author—writes the text

2. editor—judges quality and relevance

3. publisher—announces and makes available 4. archivist—preserves

5. readers—read, use, cite

+ +

Value of Journal Publication

1. certification as a significant contribution

2. standardization for discussion, citation

3. distribution to current readers

4. archiving for future readers

+ +

The Chicago Journal of Computer Science

• Published by MIT Press

• Cover theoretical computer science

• LaTEX format—already used by authors

• Market to libraries—cheaper than printed journals

• Allow copying, printing, integration in local database, …

• Distribute by ftp, Gopher, WAIS, …